OpenStudy (anonymous):

Two cars leave towns 800 kilometers apart at the same time and travel toward each other. One car's rate is 20 kilometers per hour more than the other's. If they meet in 4 hours, what is the rate of the faster car? Do not do any rounding. @Ahaanomegas

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

I showed you how to do the other one. It's similar. What is your attempt?

OpenStudy (anonymous):

i have no clue ive tried i got 80 @Ahaanomegas

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

let the speed(rate) of slower car be \( x\) therefore, speed of faster car =\(x+20\) \(\large\tt \begin{align} \color{black}{speed =\dfrac{distance}{time}\\~\\ x+x+20=\dfrac{800}{4}\\~\\ 2x+20=200\\~\\ \Large x=90\\}\end{align}\) speed of faster car = \(\Large 90+20=110\)